OpenStudy (anonymous):

what is the factored form of the expression 6x^2+13x+ 5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you

OpenStudy (solomonzelman):

yes, and the b²-4(a)(c) (the piece that is in the square root, in a quadratic formula) has to be a perfect square.

OpenStudy (anonymous):

yes it can be (3x-1) and (2x+5)

OpenStudy (aum):

6x^2+13x+ 5 = 6x^2 + 10x + 3x + 5 = 2x(3x + 5) + 1(3x + 5) = (2x + 1)(3x + 5)

OpenStudy (aum):

The above is the AC method of factoring. 6x^2+13x+ 5 In AC method you find the product of the coefficients A and C A = 6, C = 5, A*C = 30 split the middle coefficient 13 into two parts so that the sum is 13 and the product is 30. 10 + 3 = 13 10 x 3 = 30 Therefore, 6x^2 + 13x + 5 = 6x^2 + 10x + 3x + 5 and then you factor as shown in previous reply.